# BEAM: Bridging Physically-based Rendering and Gaussian Modeling for Relightable Volumetric Video

### v1

@@
Abstract
Volumetric video enables immersive experiences by capturing dynamic 3D scenes, enabling diverse applications for virtual reality, education, and telepresence. However, traditional methods struggle with fixed lighting conditions, while neural approaches face trade-offs in efficiency, quality, or adaptability for relightable scenarios. To address these limitations, we present BEAM, a novel pipeline that bridges 4D Gaussian representations with physically-based rendering (PBR) to produce high-quality, relightable volumetric videos from multi-view RGB footage. BEAM recovers detailed geometry and PBR properties via a series of available Gaussian-based techniques. It first combines Gaussian-based performance tracking with geometry-aware rasterization in a coarse-to-fine optimization framework to recover spatially and temporally consistent geometries. We further enhance Gaussian attributes by incorporating PBR properties step by step. We generate roughness via a multi-view-conditioned diffusion model, and then derive AO and base color using a 2D-to-3D strategy, incorporating a tailored Gaussian-based ray tracer for efficient visibility computation. Once recovered, these dynamic, relightable assets integrate seamlessly into traditional CG pipelines, supporting real-time rendering with deferred shading and offline rendering with ray tracing. By offering realistic, lifelike visualizations under diverse lighting conditions, BEAM opens new possibilities for interactive entertainment, storytelling, and creative visualization.
Figure 1: We present BEAM, a novel pipeline that bridges 4D Gaussians with accurate physically-based rendering to produce relightable volumetric videos, delivering immersive and realistic experiences on platforms such as VR.
[…]
1 Introduction
@@
The prevailing workflow [7, 14, 36, 19] for producing relightable volumetric videos in the industry still relies on tracked mesh sequences and texture videos, which can be seamlessly integrated into standard CG pipelines to support relighting under various lighting conditions. However, the intricate reconstruction process often introduces artifacts such as holes and noise, and the quality of relighting remains constrained, frequently resulting in visible imperfections. Neural advancements [56, 81, 74] focus on enabling relighting capabilities using neural factorization within implicit MLPs representations. However, these approaches often face challenges in balancing training efficiency, rendering speed, and output quality, ultimately failing to deliver satisfactory results. Recently, 3D Gaussian Splatting (3DGS)[35], an efficient point-based representation, has achieved photo-realistic rendering at unprecedented frame rates. While dynamic variants[30, 38, 65] can produce high-quality volumetric videos, they fail to produce the detailed geometry necessary for essential operations like relighting. Although efforts [18, 54, 40, 32] have been made to integrate physically-based rendering into the 3DGS pipeline, these methods are often computationally expensive and limited to static scenarios. These limitations severely restrict their applicability in industrial workflows, hindering the efficient production of 4D content.
@@
In this paper, we introduce BEAM, a novel pipeline that bridges 4D Gaussians with accurate physically-based rendering (PBR) for producing relightable volumetric videos from multi-view RGB footage. Our key idea is to robustly recover detailed geometry and decouple the PBR properties (e.g., ambient occlusion, roughness, and base color) using a carefully selected suite of techniques, i.e., rasterization, performance tracking, and ray tracing, all within a Gaussian-based paradigm. As a result, BEAM enables lifelike dynamic scenes that can be seamlessly and CG-friendly integrated into various platforms under diverse lightings (see Fig. 1).
@@
We first recover detailed and spatial-temporally consistent geometries from multiview video input, which organically combines the Gaussian-based performance tracking [30] with the geometry-aware Gaussian rasterization [78]. While the former excels at motion tracking and the latter at static geometry recovery, we unify them in a coarse-to-fine optimization framework. Specifically, we employ coarse joint Gaussians to track non-rigid motion and dense skin Gaussians to preserve intricate geometry details. We adopt a robust optimization process that integrates normal consistency, photometric consistency, and temporal regularization to enhance geometric accuracy and smoothness. This enables accurate depth and normal recovery from the dense Gaussians using the geometry-aware rasterizer [78], providing a robust foundation for material decomposition and relighting.
@@
We further decouple the dense 4D Gaussians to recover detailed material properties, enabling high-quality physically-based rendering grounded in the rendering equation [33] and simplified Disney BRDF [2]. Assuming human-centric scenes with negligible metallic components, we focus on accurately associating roughness, ambient occlusion (AO), and base color properties with the Gaussians, ensuring realistic and adaptable rendering under diverse lighting conditions. To achieve this, we adopt a step-by-step approach to disentangle these properties. Specifically, we first generate a roughness texture using the material diffusion module in previous work [79] with multi-view conditioning, which is associated with the dense Gaussians through UV projection. Then, for the AO and base color, we adopt a 2D-to-3D strategy, where these attributes are estimated in the input views to bake 2D material maps, and then optimize into the corresponding dense Gaussians in the 3D space. This strategy effectively reduces noise and smooths the disentanglement to improve relighting quality. For further 2D AO and based color decomposition, the lighting environment during capturing can be estimated using off-the-shelf tool [47], while the geometry attributes and roughness are obtained in previous stages. Thus, by carefully re-examining and simplifying the rendering equation [33], we identify a critical insight: both 2D AO and base color can be accurately derived by accumulated visibility information for specific points along specific directions during ray tracing. We tailor the Gaussian-based ray tracer [45] to compute such visibility, with a novel alpha blending strategy based on the dense Gaussians. This strategy efficiently captures visibility information, forming the foundation for estimating AO and base color maps in the input viewpoints.
@@
Figure 2: We propose a novel BEAM method to produce Gaussian sequences. We first use tracking results of joint Gaussians and a normal regularizer to obtain our 4D Gaussians with consistent geometry. Then we infer roughness using a generative model and apply ray tracing to compute the 2D base color and AO maps, which are then used to optimize the corresponding Gaussian attributes. Our results can be rendered under varying lighting conditions using both real-time and offline rendering.
2 Related Work
@@
In the field of Human Modeling, numerous methods [57, 60, 82, 17] have been proposed to address these challenges. Li [37] integrates temporal denoising into non-rigid mesh template tracking to capture detailed geometry, while High-quality FVV [7] compactly represents human performance using tracked mesh sequences and textured video.
With the advancement of neural rendering techniques [44], several works have incorporated time as a latent variable to handle dynamic scenes [50, 70, 62]. Meanwhile, other approaches [49, 66, 84, 22, 58, 85, 48] leverage the human parametric model [41] to reconstruct the animatable avatar. Furthermore, some hybrid methods [75, 27, 28] combine explicit volumetric fusion with implicit neural techniques to capture more details. Additionally, several methods [64, 26, 28, 53, 55] significantly accelerate training and rendering speed, by leveraging advanced data structures such as voxel grids [16], hash tables [46], and tensor decomposition [4]. Recently, 3DGS [35] ensures both high quality and fast rendering, while dynamic variants [59, 69, 42, 24, 73, 31, 15, 72, 38] enable complex 4D scene reconstruction for advanced human modeling. DualGS [30] uses joint and skin Gaussians to capture motion and detailed appearance. However, these approaches fail to recover detailed geometry and do not support relighting.
@@
Human Relighting aims to manipulate the reflectance field of the human surface, enabling an immersive fusion with novel illumination. Conventional methods [3, 9, 10, 11, 21, 23, 67, 68] use LightStage systems to capture human reflectance characteristics, requiring costly controlled lighting and dense camera arrays that are not widely accessible. In 2D image-based relighting tasks, previous methods [34, 61]rely on convolutional networks for inference, while recent diffusion-based approaches [13, 77] leverage generative models to learn complex light interactions across large datasets. However, the lack of a 3D representation makes it challenging to maintain 3D consistency of lighting. In addition, some neural relighting methods [1, 81, 76] use NeRF [44] for 3D reconstruction of material properties and lighting effects. However, their rendering quality is inadequate and hard to integrate with traditional CG pipelines. Recent methods [18, 32, 40, 20] leverages 3DGS representation for relighting due to its ability to reconstruct fine details and interaction in CG engines. For human performance relighting, researchers [6, 71, 39, 5, 83, 43] extend mesh-based and neural relighting methods by incorporating body pose priors [41, 29]. However, avatars relying on skeletal priors struggle with complex clothing, wrinkles, and human-object interactions.
3 Method
@@
Given multi-view video inputs, we aim to produce relightable 4D Gaussian sequences with physically-based rendering (PBR) materials, enabling realistic rendering under various lighting. PBR typically requires multiple material components, such as base color, metallic, roughness, normal, and ambient occlusion (AO), as shown in Fig.3. In our approach, as the metallic attribute is negligible for human bodies, we set it to zero and focus on optimizing the remaining properties: normal, roughness, AO, and base color. The complete pipeline is illustrated in Fig.2.
3.1 Gaussian Modeling and Geometry Optimization
@@
To obtain temporally consistent geometric information (depth and normal maps) for physical-based rendering, we seamlessly combine the Gaussian-based performance tracking [30] with the geometry-aware rasterizer [78] within a coarse-to-fine optimization framework. Specifically, our framework uses a dual Gaussian representation to separately model global motion and visual appearance through joint and skin Gaussians. Each skin Gaussian is anchored to multiple joint Gaussians and is warped across frames based on the tracking results of these joint Gaussians. The optimization process integrates a photometric loss $E_{\text{color}}$, a rigid term $E_{\text{rigid}}$, and a temporal regularization term $E_{\text{temp}}$.
@@
$E_{\text{normal}}=\sum_{i}\omega_{i}(1-N_{r}^{\top}N_{d}),$ (1)
@@
The total energy term in our 4D Gaussian modeling and geometric optimization framework is expressed as:
$E=\lambda_{\text{color}}E_{\text{color}}+\lambda_{\text{smooth}}E_{\text{smooth}}+\lambda_{\text{temp}}E_{\text{temp}}+\lambda_{\text{normal}}E_{\text{normal}},$ (2)
During dynamic training, $E_{\text{norm}}$ is introduced after the appearance optimization over 7,000 iterations, followed by an additional 5,000 iterations dedicated to optimizing the normals. For other energy terms not discussed here, please refer to DualGS [30]. The hyperparameters are set as follows: $\lambda_{\text{color}}=1$, $\lambda_{\text{smooth}}=0.001$, $\lambda_{\text{temp}}=0.00005$, $\lambda_{\text{normal}}=0.03$.
Figure 3: Illustration of Gaussian properties. We showcase the normal, roughness, AO and base color maps required for relighting, which are rasterized from our relightable 4D Gaussians.
3.2 PBR Materials Decomposition
To enable physical-based rendering within the Gaussian-based paradigm, we extend 3D Gaussian attributes by incorporating a roughness value and two additional third-order spherical harmonic (SH) properties for AO and base color. And we utilize a step-by-step approach to disentangle these properties.

### v2

@@
Abstract.
Volumetric video enables immersive experiences by capturing dynamic 3D scenes, enabling diverse applications for virtual reality, education, and telepresence. However, traditional methods struggle with fixed lighting conditions, while neural approaches face trade-offs in efficiency, quality, or adaptability for relightable scenarios. To address these limitations, we present BEAM, a novel pipeline that bridges 4D Gaussian representations with physically-based rendering (PBR) to produce high-quality, relightable volumetric videos from multi-view RGB footage. BEAM recovers detailed geometry and PBR properties via a series of available Gaussian-based techniques. It first combines Gaussian-based human performance tracking with geometry-aware rasterization in a coarse-to-fine optimization framework to recover spatially and temporally consistent geometries. We further enhance Gaussian attributes by incorporating PBR properties step by step. We generate roughness via a multi-view-conditioned diffusion model, and then derive AO and base color using a 2D-to-3D strategy, incorporating a tailored Gaussian-based ray tracer for efficient visibility computation. Once recovered, these dynamic, relightable assets integrate seamlessly into traditional CG pipelines, supporting real-time rendering with deferred shading and offline rendering with ray tracing. By offering realistic, lifelike visualizations under diverse lighting conditions, BEAM opens new possibilities for interactive entertainment, storytelling, and creative visualization.
Relighting, Physically-based Rendering, Human Performance Modeling, Dynamic Gaussian Splatting
journalyear: 2025copyright: ccconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755233isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Rendering
Figure 1. We present BEAM, a novel pipeline that bridges 4D Gaussians with accurate physically-based rendering to produce relightable volumetric videos, delivering immersive and realistic experiences on platforms such as VR.
1. Introduction
@@
The prevailing workflow (Collet et al., 2015; Dou et al., 2016; Lawrence et al., 2024; Garon et al., 2016) for producing relightable volumetric videos in the industry still relies on tracked mesh sequences and texture videos, which can be seamlessly integrated into standard CG pipelines to support relighting under various lighting conditions. However, the intricate reconstruction process often introduces artifacts such as holes and noise, and the quality of relighting remains constrained, frequently resulting in visible imperfections. Neural advancements (Srinivasan et al., 2021; Zhang et al., 2021; Yao et al., 2022) focus on enabling relighting capabilities using neural factorization within implicit MLPs representations. However, these approaches often face challenges in balancing training efficiency, rendering speed, and output quality, ultimately failing to deliver satisfactory results. Recently, 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023), an efficient point-based representation, has achieved photo-realistic rendering at unprecedented frame rates. While dynamic variants (Jiang et al., 2024b; Li et al., 2024a; Wang et al., 2024) can produce high-quality volumetric videos, they fail to produce the detailed geometry necessary for essential operations like relighting. Although efforts (Gao et al., 2025; Shi et al., 2023; Liang et al., 2024; Jiang et al., 2024d) have been made to integrate physically-based rendering into the 3DGS pipeline, these methods are often computationally expensive and limited to static scenarios. These limitations severely restrict their applicability in industrial workflows, hindering the efficient production of 4D content.
@@
In this paper, we introduce BEAM, a novel pipeline that bridges 4D Gaussians with accurate physically-based rendering (PBR) for producing relightable volumetric videos from multi-view RGB footage. Our key idea is to robustly recover detailed geometry of human performances and decouple the PBR properties (e.g., ambient occlusion, roughness, and base color) using a carefully selected suite of techniques, i.e., rasterization, performance tracking, and ray tracing, all within a Gaussian-based paradigm. As a result, BEAM enables lifelike dynamic scenes that can be seamlessly and CG-friendly integrated into various platforms under diverse lighting (see Fig. 1).
@@
We first recover detailed and spatial-temporally consistent geometries from multi-view video input, which organically combines the Gaussian-based performance tracking (Jiang et al., 2024b) with the geometry-aware Gaussian rasterization (Zhang et al., 2024a). While the former excels at motion tracking and the latter at static geometry recovery, we unify them in a coarse-to-fine optimization framework. Specifically, we employ coarse joint Gaussians to track non-rigid motion and dense skin Gaussians to preserve intricate geometry details. We adopt a robust optimization process that integrates normal consistency, photometric consistency, and temporal regularization to enhance geometric accuracy and smoothness. This enables accurate depth and normal recovery from the dense Gaussians using the geometry-aware rasterizer (Zhang et al., 2024a), providing a robust foundation for material decomposition and relighting.
@@
We further decouple the dense 4D Gaussians to recover detailed material properties, enabling high-quality physically-based rendering grounded in the rendering equation (Kajiya, 1986) and simplified Disney BRDF (Burley and Studios, 2012). Assuming human-centric scenes with negligible metallic components, we focus on accurately associating roughness, ambient occlusion (AO), and base color properties with the Gaussians, ensuring realistic and adaptable rendering under diverse lighting conditions. To achieve this, we adopt a step-by-step approach to disentangle these properties. Specifically, we first generate a roughness texture using the material diffusion module in previous work (Zhang et al., 2024b) with multi-view conditioning, which is associated with the dense Gaussians through UV projection. Then, for the AO and base color, we adopt a 2D-to-3D strategy, where these attributes are estimated in the input views to bake 2D material maps, and then optimized into the corresponding dense Gaussians in the 3D space. This strategy effectively reduces noise and smooths the disentanglement to improve relighting quality. For further 2D AO and base color decomposition, the lighting environment during capturing can be estimated using an off-the-shelf tool (New House Internet Services B.V., 2025), while the geometry attributes and roughness are obtained in previous stages. Thus, by carefully re-examining and simplifying the rendering equation (Kajiya, 1986), we identify a critical insight: both 2D AO and base color can be accurately derived by accumulated visibility information for specific points along specific directions during ray tracing. We tailor the Gaussian-based ray tracer (Moenne-Loccoz et al., 2024) to compute such visibility, with a novel alpha blending strategy based on the dense Gaussians. This strategy efficiently captures visibility information, forming the foundation for estimating AO and base color maps in the input viewpoints.
Figure 2. We propose a novel BEAM method to produce Gaussian sequences. We first use tracking results of joint Gaussians and a normal regularizer to obtain our 4D Gaussians with consistent geometry. Then we infer roughness using a generative model and apply ray tracing to compute the 2D base color and AO maps, which are then used to optimize the corresponding Gaussian attributes. Our results can be rendered under varying lighting conditions using both real-time and offline rendering.
@@
2. Related Work
@@
In the field of Human Modeling, numerous methods (Sun et al., 2021; Suo et al., 2021; Fridovich-Keil et al., 2023; Hu et al., 2025a) have been proposed to address these challenges. Li (Li et al., 2012) integrates temporal denoising into non-rigid mesh template tracking to capture detailed geometry, while High-quality FVV (Collet et al., 2015) compactly represents human performance using tracked mesh sequences and textured video.
Building on DynamicFusion (Newcombe et al., 2015), subsequent methods have integrated techniques like SIFT features (Innmann et al., 2016), motion field constraints (Slavcheva et al., 2017, 2018), neural deformation graphs (Bozic et al., 2021), and parametric human models like SMPL (Loper et al., 2023) to address challenges in human reconstruction, including drift and human-object interactions.
With the advancement of neural rendering techniques (Mildenhall et al., 2020), several works have incorporated time as a latent variable to handle dynamic scenes (Pumarola et al., 2021; Xian et al., 2021; Tretschk et al., 2021). Meanwhile, other approaches (Peng et al., 2021; Zheng et al., 2023; Habermann et al., 2023; Sun et al., 2025; Zhu et al., 2024) leverage the human parametric model (Loper et al., 2023) to reconstruct the animatable avatar. Furthermore, some hybrid methods (Yu et al., 2021; Jiang et al., 2022, 2023) combine explicit volumetric fusion with implicit neural techniques to capture more details. Additionally, several methods (Wang et al., 2023; Işık et al., 2023; Shao et al., 2023; Hu et al., 2025c) significantly accelerate training and rendering speed, by leveraging advanced data structures such as voxel grids (Fridovich-Keil et al., 2022), hash tables (Müller et al., 2022), and tensor decomposition (Chen et al., 2022). Recently, 3DGS (Kerbl et al., 2023) ensures both high quality and fast rendering, while dynamic variants (Sun et al., 2024; Luiten et al., 2024; Huang et al., 2024; Yang et al., 2024; Jiang et al., 2024c; Duan et al., 2024; Hu et al., 2025b; Li et al., 2024a; Jiang et al., 2025) enable complex 4D scene reconstruction for advanced human modeling. DualGS (Jiang et al., 2024b) uses joint and skin Gaussians to capture motion and detailed appearance. However, these approaches fail to recover detailed geometry and do not support relighting.
@@
Human Relighting aims to manipulate the reflectance field of the human surface, enabling an immersive fusion with novel illumination. Conventional methods (Chabert et al., 2006; Debevec, 2012; Guo et al., 2019; Hawkins et al., 2001; Wenger et al., 2005; Weyrich et al., 2006) use LightStage systems to capture human reflectance characteristics, requiring costly controlled lighting and dense camera arrays that are not widely accessible. In 2D image-based relighting tasks, previous methods (Kanamori and Endo, 2018; Tajima et al., 2021) rely on convolutional networks for inference, while recent diffusion-based approaches (Ding et al., 2023; Zeng et al., 2024) leverage generative models to learn complex light interactions across large datasets. However, the lack of a 3D representation makes it challenging to maintain 3D consistency of lighting. In addition, some neural relighting methods (Boss et al., 2021; Zhang et al., 2021; Zeng et al., 2023) use NeRF (Mildenhall et al., 2020) for 3D reconstruction of material properties and lighting effects. However, their rendering quality is inadequate and hard to integrate with traditional CG pipelines. Recent methods (Gao et al., 2025; Jiang et al., 2024d; Liang et al., 2024; Gu et al., 2024) leverages 3DGS representation for relighting due to its ability to reconstruct fine details and interaction in CG engines. For human performance relighting, researchers (Chen and Liu, 2022; Xu et al., 2024; Li et al., 2024b; Chen et al., 2024; Zheng et al., 2025; Carbonera Luvizon et al., 2024) extend mesh-based and neural relighting methods by incorporating body pose priors (Loper et al., 2023; Jiang et al., 2024a). However, avatars relying on skeletal priors struggle with complex clothing, wrinkles, and human-object interactions.
3. Method
@@
Given multi-view video inputs with known environment illumination, we aim to produce relightable 4D Gaussian sequences with physically-based rendering (PBR) materials, enabling realistic rendering under diverse lighting conditions. PBR typically requires multiple material components, such as base color, metallic, roughness, normal, and ambient occlusion (AO). As the metallic attribute is negligible for human bodies, we set it to zero and focus on optimizing the remaining properties: normal, roughness, AO, and base color. The resulting relightable 4D Gaussians are rendered under various lighting and scene configurations, and further integrated into CG engines and VR applications. The complete pipeline is illustrated in Fig. 2.
3.1. Gaussian Modeling and Geometry Optimization
@@
To obtain temporally consistent geometric information (depth and normal maps) for physically-based rendering, we seamlessly combine the Gaussian-based performance tracking (Jiang et al., 2024b) with the geometry-aware rasterizer (Zhang et al., 2024a) within a coarse-to-fine optimization framework. Specifically, our framework uses a dual Gaussian representation to separately model global motion and visual appearance through joint and skin Gaussians. Each skin Gaussian is anchored to multiple joint Gaussians and is warped across frames based on the tracking results of these joint Gaussians. The optimization process integrates a photometric loss $E_{\text{color}}$, a smooth term $E_{\text{smooth}}$, and a temporal regularization term $E_{\text{temp}}$.
@@
(1) $E_{\text{normal}}=\sum_{i}\omega_{i}(1-N_{r}^{\top}N_{d}),$
@@
Together, the total energy term in our 4D Gaussian modeling and geometric optimization framework is expressed as:
(2) $E=\lambda_{\text{color}}E_{\text{color}}+\lambda_{\text{smooth}}E_{\text{smooth}}+\lambda_{\text{temp}}E_{\text{temp}}+\lambda_{\text{normal}}E_{\text{normal}},$
During dynamic training, $E_{\text{norm}}$ is introduced after the appearance optimization over 7,000 iterations, followed by an additional 5,000 iterations dedicated to optimizing the normals. For other energy terms not discussed here, please refer to DualGS (Jiang et al., 2024b). The hyperparameters are set as follows: $\lambda_{\text{color}}=1$, $\lambda_{\text{smooth}}=0.001$, $\lambda_{\text{temp}}=0.00005$, $\lambda_{\text{normal}}=0.03$.
With the above energy term, we obtain a Gaussian sequence with temporally consistent geometry. For further mesh extraction, we render depth maps for training views and then adopt TSDF (Curless and Levoy, 1996) fusion. The resulting mesh sequence is then used for subsequent PBR material decomposition.
Figure 3. Gallery of our results. We present some real-time rendering results under HDRI settings, which deliver high-fidelity rendering of human performances across challenging motions and complex clothing textures.
3.2. PBR Materials Decomposition
To enable physically-based rendering within the Gaussian-based paradigm, we enhance the 3D Gaussian representation by introducing additional material attributes: roughness $r\in[0,1]$, ambient occlusion (AO) $\mathcal{A}\in[0,1]$, and base color $\boldsymbol{\rho}\in[0,1]^{3}$. The $i^{th}$ extended Gaussian $\mathcal{P}_{i}(t)$ at time $t$ is defined as:
(3) $\mathcal{P}_{i}(t)=\{\boldsymbol{\mu}_{i}(t),\boldsymbol{q}_{i}(t),\boldsymbol{s}_{i}(t),o_{i}(t),\boldsymbol{c}_{i}(t),\boldsymbol{\rho}_{i}(t),r_{i}(t),\mathcal{A}_{i}(t)\},$
where each component denotes position, orientation, scale, opacity, color, base color, roughness, and ambient occlusion, respectively. To better capture subtle appearance changes over time, we allow the base color $\rho$ to vary for each Gaussian. The roughness $r$, on the other hand, is kept time-invariant as a stable material property. To disentangle these material properties, we adopt a step-by-step optimization strategy.